OpenStudy (anonymous):

3P-5=3(P+1) A) 0 only B) no Solution C)3 only D)all real Numbers PLZ help

OpenStudy (amoodarya):

3p-5=3(p+1) 3p-5=3p+3 3p emitted from both side so -5=3 that is impossible so the equation has no root

OpenStudy (anonymous):

thx